OpenStudy (anonymous):

Use the quadratic formula to solve the equation: 3x^2-7x=5 use radicals as needed, express complex numbers in terms of i

OpenStudy (anonymous):

do you know delta operator ?

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

ok i will do without ^^

OpenStudy (anonymous):

first step is : \[3 * (x^{2} - 7 / 3*x -5/3) = 0\]

OpenStudy (anonymous):

after that you have to do a little trick, it is based on : (x - a)^2 = x^2 -2*a*x + a^2 do you know the trick ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

okay, so you have to make the x^2 disappear because we don't know what to do with it

OpenStudy (anonymous):

\[3*( (x-7/6)^{2} -5/3 - (7/6)^{2})=0\] is the way to do it

OpenStudy (anonymous):

because (x - 7/6)^2 = x^2 - 7/3x + 7/6 ^2

OpenStudy (anonymous):

first step=subtract 5 from both sides

OpenStudy (radar):

The problem specifically requires you use the quadratic formula, I suggest you do so. The quadratic formula is:\[-b \pm \sqrt{b ^{2}-4ac}\over2a\] where a is the coefficient of x^2, b is coefficient of x and c is the constant, the the quadratic is placed in standars form:\[ax ^{2}+bx+c=0\] Have you used the quadratic formula before?

OpenStudy (anonymous):

dont see where the 'i' would come into place since youre not taking the root of a -

OpenStudy (anonymous):

if u graph it does intersect with the x-ax

OpenStudy (anonymous):

i think there are plenty of those exercise and the text is the same for the other

OpenStudy (anonymous):

no c= -5

OpenStudy (anonymous):

yes but she said she didn't jnow the delta operator :s

OpenStudy (radar):

\[3x ^{2}-7x=5\] Placing it in standard form\[3x ^{2}-7x-5=0\]a=3, b=-7, and c=-5 Just plug these values in the formula and solve.

OpenStudy (anonymous):

\[\frac{7\pm\sqrt{7^2-4\times(3)\times5}}{2\times3}\]

OpenStudy (anonymous):

so wouldnt it be -7 +or- the sqrt of 109. take ans and divide by 6

OpenStudy (radar):

There it is, with the numbers plugged in. What more could you ask for?

OpenStudy (anonymous):

to understand ?

OpenStudy (anonymous):

yea just plug in accordingly to formula

OpenStudy (anonymous):

aha it's so meaningless : use the formula T_T

OpenStudy (radar):

That was the purpose of the problem, learn to read and do.

OpenStudy (radar):

Look at the top of this page and you will see her work instructed to use the quadratic formula, not competeing the square, or factoring.

OpenStudy (anonymous):

yes, but so, why don't she knows the delta operator ? you don't learn it ?

OpenStudy (anonymous):

somehow site is very slow. i had a typo but you are right, it is \[x=\frac{7\pm\sqrt{109}}{6}\]sorry it took a while delta operator?

OpenStudy (anonymous):

ow okay hum sorry for that, delta is b^2 - 4*a*c in france :s because it's use for two variables equations as well ^^"

OpenStudy (anonymous):

well i guess we used it right?

OpenStudy (radar):

Is that what we call the discriminate?

OpenStudy (anonymous):

yep, but i asked her if she knew it :) she didn't because it hasn t a name for her, so i thought she wanted the calculus

OpenStudy (anonymous):

yes it is discriminant or delta in french

OpenStudy (radar):

She has moved on. So maybe later.

OpenStudy (anonymous):

Wow, what a mess under this problem. It seems those with less experience who are providing me answers today are providing me wrong answers :( Perhaps I'll wait for the heros and legends and such from now on...

OpenStudy (anonymous):

all the answer are gool ;) mine is without the formula, so it's not what you wanted sorry ;), the one of satellite is a good one. (else you can learn your formula, it's an exercise to see if you know your lesson :) )

OpenStudy (anonymous):

good *